OpenStudy (mysticfury):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 3 –2 4 Line 2 x y 0 8 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (mysticfury):

@welshfella @JDcap , i chose B again

OpenStudy (jdcap):

y=x/2+5 y=8-x x=2, y=6

OpenStudy (mysticfury):

so is my answer correct then? @JDcap

OpenStudy (jdcap):

Yes

OpenStudy (mysticfury):

okay i got one more